OpenStudy (anonymous):

Solve 2x2 – 8x = –7.

OpenStudy (anonymous):

x = 1/2 (4-sqrt(2)) x = 1/2 (4+sqrt(2)) Use the quadratic formula

OpenStudy (anonymous):

possible answers are -2+-sqrt2, -1+-sqrt2, 2+-sqrt2/2, 2+-sqrt2/2